OpenStudy (anonymous):

In an experiment to find the formula of the oxide formed of the element M, 5.5g of M was burnt in oxygen. The mass of the oxide was 8.7g What is the formula of the metal oxide?

OpenStudy (anonymous):

@Callisto sorry...

sam (.sam.):

\[M+O_2--->MO\] M=5.5g MO=8.7g \[8.7g-5.5g=3.2g O_2\]

OpenStudy (anonymous):

I got MO2

sam (.sam.):

So which metal have you predicted?

OpenStudy (anonymous):

they are just asking the formula, with "M"

OpenStudy (callisto):

If it is MO2, then it's probably a group II metal... but I can't find it!!! (Sorry.. starving and must eat something now...)

sam (.sam.):

If metal + oxygen , it must be 2MO or M2O otherwise the charges wouldn't cancel \[2M+O2−−−>2MO\] \[4M+O2−−−>2M2O\] Unless it is transition metal

OpenStudy (callisto):

Why can't it be M2O3? Like Al2O3 (even though it doesn't seem like to be the answer :|)

OpenStudy (callisto):

So... is RAM of M = 55?

OpenStudy (callisto):

RAM = relative atomic mass I think you missed some data given in the question, right?

OpenStudy (callisto):

@zaphod

OpenStudy (kayne):

Firstly, what's the equation to solve this problem? Is it: xM+(y/2)O2---->MxOy...??

OpenStudy (callisto):

It's difficult to solve that in that way.... :( Perhaps insufficient information.

OpenStudy (kayne):

Yups.. I agree @Callisto

OpenStudy (callisto):

If in that way, no. of mole of O2 used = (8.7-5.5)/(16.0x2) = 0.1mole y/2 = 0.1 => y=0.2 For no. of mole of M: 5.5/m = x For no. of mole of MxOy: 8.7 / (xm+16.0y) = 1 8.7 / (5.5 + 16.0x0.2) =1 1=1 ._.

OpenStudy (callisto):

If the question is this, then it's easy ._. In an experiment to find the formula of the oxide formed of the element M, 5.5 g of M was burnt in oxygen. The mass of the oxide was 8.7 g. [Ar: M, 55; O, 16.] What is the formula of the metal oxide? Then, no of mole of M = 5.5/55 = 0.1mole no. of mole of O2 = (8.7-5.5) / (16.0x2) = 3.2/32 = 0.1 moles Since the mole ratio of M:O2 = 0.1/0.1 = 1:1 So the formula of metal oxide is MO

OpenStudy (callisto):

Note: I just googled this question and find that it's one of the question in UNIVERSITY OF CAMBRIDGE INTERNATIONAL EXAMINATIONS - General Certificate of Education Ordinary Level (October/November 2004) It's question 7

OpenStudy (anonymous):

but the answer is MO2, give me the link

OpenStudy (anonymous):

yeah but the answer says its MO2

OpenStudy (callisto):

:O Wait... let me check my answer again...Sorry!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

it should be O ----> not O2

OpenStudy (callisto):

Wait?! What do you mean?

OpenStudy (callisto):

The stupid problem of mine I've found is that.... Ratio of M:O2 = 1:1 So, the formula is MO2 ._. M + O2 -> MO2

OpenStudy (anonymous):

the oxygen M + O ----> 5.5/55 + 3.2/16 ---> 0.1 + 0.2 0.2/0.1 --->2 Oxygen MO2.......AM i right

OpenStudy (callisto):

No... oxygen gas is O2...

OpenStudy (kayne):

Hey I got the MO2.. Wait.. I'll post my working..

OpenStudy (anonymous):

okay..

OpenStudy (callisto):

no of mole of M = 5.5/55 = 0.1mole no. of mole of O2 = (8.7-5.5) / (16.0x2) = 3.2/32 = 0.1 moles Since the mole ratio of M:O2 = 0.1/0.1 = 1:1 So the formula of metal oxide is MO2 M + O2 -> MO2 My bad :|

OpenStudy (callisto):